# The Schwurbelarchiv: a German Language Telegram dataset for the Study of Conspiracy Theories

### v2

@@ 4.6 Transcription
At the time of this publication, transcription of all files with audio tracks is ongoing. New files will be iteratively added to the online dataset. For the transcription process, we used the openai/whisper-large-v3-turbo model, which was the latest whisper model available at the time of publication Radford et al. (2022). To boost the transcription speed, we combined the model with FlashAttention-2 Dao (2024) which resulted in a transcription rate of around 500x realtime. This increases the amount of text in our dataset from 1,490M words in the written messages to a total of approximately 2,103M words (estimation of video file words, see Table 4) the text corpus by 41%. As media files comprise a considerable share of all forwarded messages, transcribed messages are an important carrier of information between chats. A feature of the text generated via transcription is that it consists of a well-defined vocabulary and therefore has no typos or alternative spellings. While this precludes the analysis of textual characteristics related to these features, it might allow for more accurate text analysis and topic identification and requires less pre-processing before further analysis.
Audio Video Images
Count 0.889M 2.475M 11.71M
Total words 217.8M 395.1M*
Total filesize 15.1TB 20.2TB 1.14TB
Total duration 63,395 h 63,194 h
Median duration 74.96 s 42.57 s
Table 4: Statistics of transcribable media files. Video word count was estimated from a random sample of 6.2K videos.
Next to voice messages and video files, 77% of the 15,065,197 media files that were sent by Telegram users in the dataset were images. Manual investigation of a small fraction of the images suggests that many of them contain text in form of memes, slogans, or announcements. While we believe it technically feasible to extract text from these images or even use new multimodal large language models to create text descriptions of them, such processing incurs additional computational processing cost.
@@ 5 Discussion and Conclusions
In this paper, we present the Schwurbelarchiv dataset, a large-scale collection of more than 6,000 Telegram chats collected via snowball-sampling, focusing primarily on German-language discussions in the realm of conspiracy theories, misinformation, and political extremism in the time from 2016 to 2022. Notably, posting activity in the dataset peaks during the COVID-19 pandemic and at key political events of the time, indicating heightened and potentially coordinated online activity in these special interest groups.
With the present data publication, we provide a cleaned and validated version of the dataset that is easy to access and re-use for other researchers. We clarify questions around the ethics of using the dataset as well as the terms of use. The dataset contains digital trace data of individuals. While we pseudonymised usernames, the dataset is not completely anonymised and in principle allows for the re-identification of individuals based on publicly disclosed information contained in the data. We release this dataset because we believe the benefit of being able to study the spread of conspiracy theories in a completely unregulated online environment at scale during a time of crisis outweighs the risk of infringing on user expectations and privacy. However, we do not intend this dataset to be used for the profiling of individuals and strongly caution against any attempts of re-identification.
Rather than relying on data collection from the Telegram API, the dataset was scraped with an automated approach via Telegram desktop clients that enabled the download of multimedia content, e.g. images, videos and voice messages. While the raw dataset was accessible in the Internet Archive before, we make it easier to use for research by providing an assessment of its completeness and scope, parsing and cleaning it. Furthermore and next to the large corpus of text messages, we provide transcriptions of voice messages and videos, enriching the corpus with more than three million files which we make accessible for further analysis. The Schwurbelarchiv is – to our knowledge – the only large-scale Telegram dataset that includes transcribed audio content. Based on our qualitative observations of the chats in the dataset, the discussion is often driven by video or voice message contributions while text messages are mostly used to comment on this multimedia content. Therefore, by ignoring multimedia content – as is often-times the case in research that uses digital trace data – large parts of the discussion are missed. With our work we therefore aim to enable research into the role of multimodal content in spreading conspiratorial narratives.
@@
We assessed the completeness of our dataset by comparing it to the content of comparable Telegram datasets Baumgartner et al. (2020); Mohr (2023); Zehring and Domahidi (2023). This comparison also serves to estimate the scope of conspiracy-related discourse on Telegram during the COVID-19 pandemic in German speaking countries. On the chat level the overlap with other datasets is low, with a maximum of 13% of the chats shared with Mohr (2023). This low overlap can largely be explained by a difference in inclusion criteria used to select chats in the different dataset, and differences in the temporal scope of the data sets. Another way to gauge the completeness of our dataset regarding conspiracy-related discourse is to look at message forwards. Here, we found that 47.9% of incoming forwarded messages are also contained as original messages in our dataset. On the flip side, this means that around half of the forwarded messages originate in chats that are not contained in our dataset. We therefore estimate that we cover around half of the conspiracy-related discourse happening in predominantly German-language Telegram chats at the time in our dataset. As we have no reason to believe that the selection of chats included in our dataset is systematically biased in any way, other than being centred on conspiracy-related discourse in German, we believe that the Schwurbelarchiv is an adequate resource for the study of conspiracy theories in the given temporal and geographic scope.
@@
That being said, another limitation of our dataset relates to the unknown point in time at which data was collected for each chat. As has been reported by Buehling (2024), messages on Telegram are deleted relatively frequently, and deletion patterns are not random. We estimate large portions of our dataset, particularly towards the end of the observation period, to have been collected via streaming and therefore be relatively unaffected by message deletion. However, we cannot determine the point at which streaming started for each chat. Nevertheless, we note that the other datasets covering a similar geographical and temporal scope have all been collected by bulk download through the Telegram API Zehring and Domahidi (2023); Mohr (2023); Baumgartner et al. (2020). Therefore, these data sets are subject to increasing message deletion the farther a point in time is away from the time of data collection as well. By including both bulk-downloaded and streamed portions of chat content, the Schwurbelarchiv is likely more complete in terms of deleted messages than other data sets.
@@
Next to the content of the messages, our data offers insights into the dynamics of information sharing and highlights how message frequency, group activity, and author engagement are distributed and evolve over time. The activity of authors in groups exhibits heavy-tailed distributions that approximate a power-law. This is consistent with activity patterns across social media platforms, such as Twitter Orellana-Rodriguez and Keane (2018). Furthermore, activity distribution is very uneven, with 80% of messages being posted in the top 12.1% of chats 80% of messages being posted by the top 3% of authors.

### v3

@@ 4.6 Transcription
Uniquely, the dataset contains transcriptions of both audio messages and video files. For the transcription process, we used the openai/whisper-large-v3-turbo model, which was the latest whisper model available at the time of publication [20]. This increases the amount of text in our dataset from 1,761M words in the written messages to a total of approximately 2,718M words (see Table 4), increasing the text corpus by 54.3%. As media files comprise a considerable share of all forwarded messages, transcribed messages are an important carrier of information between chats. A feature of the text generated via transcription is that it consists of a well-defined vocabulary and therefore has no typos or alternative spellings. While this precludes the analysis of textual characteristics related to these features, it might allow for more accurate text analysis and topic identification and requires less pre-processing before further analysis. Regarding transcription quality, the Whisper large-v3-turbo model used for transcription achieves Word Error Rates (WER) between approximately 4% and 7% for German on standard benchmarks such as Common Voice 15 and Fleurs [20]. However, these benchmarks consist of clean, read-aloud speech and therefore represent a best-case transcription accuracy. The audio content in the Schwurbelarchiv includes potentially incomprehensible speech features like informality, videos with background noise, dialect speech, all of which are likely to increase error rates. To accommodate these challenges, we employed Voice Activity Detection (VAD) to filter non-speech segments, to help mitigate noise from non-voice signals. As for potential improvements: speaker diarization and the use of the full (non-turbo) Whisper model would have offered potential quality improvements but were out of scope due to the substantially higher computational cost. We consider the chosen pipeline to represent the best trade-off between transcription quality and feasibility at this scale. We did not conduct a manual evaluation of transcription quality on a sample of our data, which constitutes a limitation of this work.
Text Voice messages Video Images
Count 54.10M 0.87M 1.58M 11.87M
Total words 1,761M 530M 427M
Total filesize 0.91 TB 17.43 TB 1.14 TB
Total duration 65,433 h 61,406 h
Median duration 76.83 s 74.00 s
Table 4: Multimodal data statistics. Excludes video and audio files that were corrupted, didn’t include an audio track or had an unknown codec.
Next to voice messages and video files, 78% of the 15.3M media files that were sent by Telegram users in the dataset were images. Manual investigation of a small fraction of the images suggests that many of them contain text in form of memes, slogans, or announcements. While we believe it technically feasible to extract text from these images or even use new multimodal large language models to create text descriptions of them, such processing incurs additional computational processing cost and is left to future work.
@@ 5 Discussion and Conclusions
In this paper, we present the Schwurbelarchiv dataset, a large-scale collection of more than 5,800 Telegram chats, focusing primarily on German-language discussions in the realm of conspiracy theories, misinformation, and political extremism in the time from 2016 to 2022. Notably, posting activity in the dataset peaks during the COVID-19 pandemic and at key political events of the time, indicating heightened and potentially coordinated online activity in these special interest groups.
With the present data publication, we provide a cleaned and validated version of the dataset that is easy to access and re-use for other researchers. We clarify questions around the ethics of using the dataset as well as the terms of use. The dataset contains digital trace data of individuals. While we pseudonymized usernames, the dataset is not completely anonymised and in principle allows for the re-identification of individuals based on publicly disclosed information contained in the data. We release this dataset because we believe the benefit of being able to study the spread of conspiracy theories in a completely unregulated online environment at scale during a time of crisis outweighs the risk of infringing on user expectations and privacy. In particular, what we add with this publication is data cleaning and validation of a data set that was originally collected by an anonymous data hoarder, which makes the data set more usable for researchers. Malicious use that attempts to single out individuals is already possible with the original data set available on the Internet Archive. However, we emphasize that we do not intend this dataset to be used for the profiling of individuals and strongly caution against any attempts of re-identification.
Rather than relying on data collection from the Telegram API, the dataset was scraped with an automated approach via Telegram desktop clients that enabled the download of multimedia content, e.g. images, videos and voice messages. While the raw dataset was accessible on the Internet Archive before, we make it easier to use for research by providing an assessment of its completeness and scope, parsing and cleaning it. Furthermore and next to the large corpus of text messages, we provide transcriptions of voice messages and videos, enriching the corpus with more than 126,000 hours of transcribed audio and video content which we make accessible for further analysis. The Schwurbelarchiv is – to our knowledge – the only large-scale Telegram dataset that includes transcribed audio content. Based on our qualitative observations of the chats in the dataset, the discussion is often driven by video or voice message contributions while text messages are mostly used to comment on this multimedia content. Therefore, by ignoring multimedia content – as is often-times the case in research that uses digital trace data – large parts of the discussion are missed. With our work we therefore aim to enable research into the role of such content in spreading conspiratorial narratives.
@@
We assessed the completeness of our dataset by comparing it to the content of comparable Telegram datasets [3, 13, 9]. This comparison also serves to estimate the scope of conspiracy-related discourse on Telegram during the COVID-19 pandemic in German speaking countries. On the level of chats the overlap with other datasets is low, with a maximum of 13% of the chats contained in the Schwurbelarchiv shared with Mohr [13]. This low overlap can largely be explained by a difference in inclusion criteria used to select chats, and differences in the temporal scope of the data sets. Another way to gauge the completeness of our dataset regarding conspiracy-related discourse is to look at message forwards. Here, we found that 54.1% of incoming forwarded messages are also contained as original messages in our dataset. On the flip side, this means that around half of the forwarded messages originate in chats that are not contained in our dataset. We therefore estimate that we cover around half of the conspiracy-related discourse happening in predominantly German-language Telegram chats at the time in our dataset. Other than potentially undersampling smaller chats from which fewer messages are forwarded, we have no reason to believe that the selection of chats included in our dataset is systematically biased. We therefore believe that the Schwurbelarchiv is an adequate resource for the study of conspiracy theories in the given temporal and geographic scope.
@@
That being said, another limitation of our dataset relates to the unknown point in time at which data was collected for each chat. As has been reported by Buehling [15], messages on Telegram are deleted relatively frequently, and deletion patterns are not random. We estimate large portions of our dataset, particularly towards the end of the observation period, to have been collected via streaming and therefore be relatively unaffected by message deletion. However, we cannot determine the point at which streaming started for each chat. Nevertheless, we note that the other datasets covering a similar geographical and temporal scope have all been collected by bulk download through the Telegram API [9, 13, 3]. Therefore, these data sets are subject to increasing message deletion the farther a point in time is away from the time of data collection as well. By including both bulk-downloaded and streamed portions of chat content, the Schwurbelarchiv is likely more complete in terms of deleted messages than other data sets.
@@
Next to the content of the messages, our data offers insights into the dynamics of information sharing and highlights how message frequency, group activity, and author engagement are distributed and evolve over time. The activity of authors in groups exhibits heavy-tailed distributions that approximate a power-law. This is consistent with activity patterns across social media platforms, such as Twitter [17]. Furthermore, activity distribution is very uneven, with the top 10% chats aggregating 73.3% of all messages and the top 10% authors contributing 94.6% of all messages.
@@
Declarations
[…]
Availability of data and materials The original Schwurbelarchiv dataset is available at the Internet Archive  [^11] https://github.com/cs2-lab/schwurbelarchiv. The cleaned and processed version of the dataset described in the paper is available upon reasonable request. Metadata of the dataset is available on Zenodo  [^12] https://zenodo.org/records/14704628. All code used to parse and analyze the data is publicly available on GitHub  [^13] https://github.com/cs2-lab/schwurbelarchiv.
[…]
Competing interests The author declare no competing interests.
[…]
Funding This research was funded in whole or in part by the Austrian Science Fund (FWF) [DOI:10.55776/P37280]. For open access purposes, the author has applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. MA, EH & JP acknowledge the FWF funding.
[…]
Authors’ contributions MA, JL & JP conceived and designed the study. MA, EH, JL & JP performed data analysis and cleaning. MA, EH, JL & JP drafted the manuscript. All authors read and approved the final manuscript.
[…]
Acknowledgments The authors thank the anonymous individual who collected the original dataset, which made this work possible.
Appendix A Column description of the dataset
Column Description
uuid unique identifier for each row
folder_id internal folder identifier for the group/channel
chat_type type of chat (channel or group) based on number of different authors within a group (>3>3 authors == group); includes manual classification for 37 chat_names
chat_name name of the group/channel
author hashed name of author
link_url link as found in HTML a-tag
media_file name of attached media file
media_file_type voice message, video, or photo
message message content
message_id ID of message in respective group
posting_date date and time of message in the group (UTC-aware timestamptz)
replied_to message_id of message that has been replied to
website website domain, if link_url refers to a website
fwd_link_url link from forwarded message
fwd_media_file internal path of the forwarded media file
fwd_media_file_type voice message, video, or photo
fwd_message forwarded message content
fwd_posting_date date and time of message in the original group
message_lang language detected in message
is_conspiracy whether message was classified as conspiracy content
audio_transcription transcribed content of a speech-containing audio file
audio_length_s duration of an audio file (in seconds)
audio_size_mb file size of the attached audio file (in MB)
audio_lang language detected in audio transcription
video_transcription transcribed content of a speech-containing video file
video_length_s duration of a video file (in seconds)
video_lang language detected in video transcription
fwd_posting_date date and time of message in the original group (UTC-aware timestamptz)
fwd_author hashed name of the original author of the forwarded message
is_duplicate boolean flag marking rows that duplicate another due to redundant scraping of that group. Lines kept and flagged instead of deleted as for some semantic duplicates, the replied_to column references to only one of these lines.
Table 5: Schwurbelarchiv dataset columns.
Appendix B Base list of English keywords by [13]
keywords = { ’covid’, ’corona’, ’virus’, ’pandemic’, ’lockdown’, ’health’, ’mask’, ’distancing’, ’outbreak’, ’symptom’, ’quarantine’, ’influenza’, ’vaccine’, ’vaccination’, ’pandemic’, ’ventilator’, ’isolation’, ’immunity’, ’hospital’, ’icu’, ’intensive care unit’, ’treatment’, ’virologist’, ’clinic’, ’homeopathy’, ’sick’, ’pharma’, ’polymerase chain reaction’, ’pcr’, ’pertussis’, ’emergency’, ’injection’, ’doctors’, ’cellular’, ’remote work’, ’frontline’, ’covid-19’, ’sars’, ’sanitation’, ’pathogen’, ’propaganda’, ’disease’, ’epidemic’, ’diarrhea’, ’adjuvants’, ’respiratory’, ’hygiene’, ’protein’, ’medicine’, ’new cases’, ’positive’, ’scientist’, ’contagious’, ’mandate’, ’variant’, ’infect’, ’deaths’, ’ill’, ’cough’, ’measure’, ’viral’, ’mrna’, ’prevent’, ’healthcare’, ’contract’, ’shutdown’, ’smallpox’, ’booster’, ’antibody’, ’dose’, ’evidence’, ’misinformation’, ’isolation’, ’observed’, ’mandatory’, ’allergic’, ’allergy’, ’immune’, ’shortage’, ’syndrome’, ’drug’, ’chinese’, ’test’, ’restriction’, ’spread’, ’vax’, ’experimental’ }